Question 3: Looking at Amazons 2017 10K filing, what is the debt to equity ratio on the balance sheet and what type of debt and equity instruments are on the balance sheet (e.g. term loans with due date PIK notes, common stock, warrants)? 43 words
Answer #1

Debt Equity Ratio = Total Liabilities / Total Equity

As on 31st Dec 2017, Amazons Total Liabilities (Long and Short Term) are = (57883 +45718) =$103601 Mn.

Stakeholders Equity = $ 27709 Mn.

Debt to Equity Ratio = 103601 / 27709  = 3.73 Times.

Type of Equity Instruments reflecting on Balance Sheet are, Common Stock and Treasury Stock.

Accumulated Other Comprehensive income and Retained earnings are not equity instruments but form part of stakehtolder equity.

Type of Debt Instruments : Long term debt contains Unsecured Senior notes of $ 24.3 Billion.

Other Long term debt is a secured Credit Facility from a lender.